OpenStudy (anonymous):

The whole thing is a fraction. K N^t =-------------- [K-N^3] 1+[------]e^-rt [ N^3 ] Solve for N ...yep.

myininaya (myininaya):

\[N^t=\frac{K}{1+\frac{K-N^3}{N^3}e^{-rt}}\]?

OpenStudy (anonymous):

That's what is is.

myininaya (myininaya):

well hmm.. I don't know if we can solve it for N but we can try - I would first get rid of that compound fraction

myininaya (myininaya):

have you done that?

myininaya (myininaya):

or know how?

OpenStudy (anonymous):

Class: 2+2 Homework: 145*3421 Quiz: 232/121*362^3

OpenStudy (anonymous):

I have no idea how to do ANY of this stuff.

myininaya (myininaya):

I'm sorry that does that mean

myininaya (myininaya):

what does that mean*

myininaya (myininaya):

2+2=4...

OpenStudy (anonymous):

It means that either we never went over this stuff, or I wasn't paying attention.

myininaya (myininaya):

i don't know what class homwork, quiz has to do with this problem

OpenStudy (anonymous):

Just...never mind. What I'm saying that in class it's easy, homework...meh, quiz=WHAT THE HECK IS THIS

myininaya (myininaya):

oh I thought you were trying to give me additional stuff to the problem

OpenStudy (anonymous):

Oh.

myininaya (myininaya):

\[N^t=\frac{K}{1+\frac{K-N^3}{N^3e^{rt}}} \cdot \frac{N^3e^{rt}}{N^3e^{rt}} \] I'm going to multiply top and bottom by N^3e^{rt} so I can get rid of the compound fraction tell me what we have as a result

OpenStudy (anonymous):

It's multiple choice, if that helps at all.

OpenStudy (anonymous):

A. N^1 +e^-rt KN^t-3 -e^-rt N^1 =K B. N^1 +K^-rt KN^t-4 -e^-rt N^1 =K C. N^1 +e^-rt KN^t-3 -e^-rt K^1 =N D. K^4 +e^-rt KK^t-3 -e^-rt N^1 =K

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

It is weird the directions say to solve for N when that isn't what happened

myininaya (myininaya):

Like there is still N on both sides of the equation on C

myininaya (myininaya):

Can you post a picture of the question?

myininaya (myininaya):

I want to see the exact wording.

myininaya (myininaya):

and see if the actual equation is right

myininaya (myininaya):

unless @mathstudent55 can use his esp skills to figure out what we are suppose to do :p

myininaya (myininaya):

I wonder why they wrote N^1 instead of N or K^1 instead of K

myininaya (myininaya):

Like I said I think I need to see a picture of this question